# Parallel transport in rotating frames and projective holonomic quantum computation

### v2

@@ Abstract
Nonadiabatic holonomic quantum computation is a promising approach for implementing quantum gates, offering both efficiency and robustness against certain types of errors. A key element of this approach is a geometric constraint known as the parallel transport condition. In line with the principle of covariance, this condition must be appropriately modified when changing reference frames. In this paper, we detail how to adjust the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame. Furthermore, building on gauge invariance considerations, we develop a framework for nonadiabatic holonomic quantum computation with projective gates. The parallel transport condition of this framework effectively addresses the issue of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. Finally, we extend the isoholonomic inequality to projective gates, establishing a minimum execution time for projective holonomic quantum gates.
@@ I   Introduction
Nonadiabatic holonomic quantum computation [1, 2, 3] has been introduced as a fast alternative to its well-established, robust, but relatively slow adiabatic counterpart [4, 5, 6]. While adiabatic holonomic quantum computation relies on the gradual manipulation of control parameters, nonadiabatic holonomic quantum computation implements quantum gates through internal dynamics driven by parallel transporting Hamiltonians. These Hamiltonians, fundamental to the geometric nature of the gates, evolve the computational subspace without inducing any time-local rotations within it, ensuring that the computation is governed purely by geometric phases rather than dynamical effects.
In the laboratory frame, the parallel transport condition can be formulated as the requirement that the velocity of any computational basis remains within the kernel bundle of the non-Abelian Aharonov-Anandan connection [7, 8]. To simplify the analysis of quantum dynamical equations, a common approach is to switch to a rotating reference frame. However, to maintain physical consistency, the dynamical equations must be transformed covariantly when changing reference frames. After outlining key concepts in holonomic quantum computing, we work out how to transform the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame.
@@
The isoholonomic inequality, a fundamental result in holonomic quantum computation [18, 19, 20], establishes a sharp lower bound on the length of the loops along which the computational subspace can be parallel transported to holonomically implement a given quantum gate [20]. We extend this inequality to projective quantum gates and, as a corollary, provide an estimate for the execution time of projective holonomic quantum gates. This extension offers deeper insight into the geometric constraints governing the implementation of quantum gates in the projective setting.
@@
In this paper we use the word frame in two different senses. First, there are frames of reference used to describe the entire quantum system, referred to as laboratory or rotating frames. These provide a basis for understanding the system’s overall behavior. Second, we introduce frames that span the computational space, which we call computational frames or $n$-frames. The context should make it clear which type of frame is being discussed.
@@ II   Conventional nonadiabatic holonomic quantum computation
Holonomic quantum computing uses holonomy to implement gates. Holonomy refers to the geometric phenomenon that when the computational space is transported along a closed path, a co-transported computational frame, that is, a frame spanning the computational space, may end up rotated relative to its original configuration, even though the transport is parallel, that is, performed without causing any time-local rotation of the computational frame.
A holonomic gate is a unitary operator representing the rotation of a computational frame resulting from a cyclic parallel transport of the computational space. A more precise description of a holonomic gate is as follows (see also [1, 2, 3, 20]). Consider a quantum system with a Hilbert space $\mathcal{H}$ and a computational subspace $\mathcal{R}$ [^1] Thus, computational input states are prepared to have support in $\mathcal{R}$, and are manipulated by unitary operators on $\mathcal{R}$, a.k.a gates. Let $\Pi_{t}$ be a one-parameter family of isometric embeddings of $\mathcal{R}$ in $\mathcal{H}$ that transport $\mathcal{R}$ in a loop:
@@
The transport is said to be parallel if, for each vector $|\psi\rangle$ in $\mathcal{R}$ and at every time $t$, the curve $|\psi_{t}\rangle=\Pi_{t}|\psi\rangle$ intersects the displaced computational space perpendicularly:
@@
Geometrically, condition (2) means that the operators $\Pi_{t}$ do not cause any time-local rotation within the computational space. The resulting rotation of a computational frame is thus solely a consequence of the translational motion of $\mathcal{R}$ through the Hilbert space. A crucial fact is that any curve of equidimensional subspaces of $\mathcal{H}$ can be realized as a transport by a unique one-parameter family of parallel transporting operators. These operators are the parallel transport operators associated with the curve. [^2] We assume that the parameter of any one-parameter family—a curve—of states, operators, or spaces ranges from 00 to $\tau$, and that the family depends piecewise smoothly on the parameter.
In practice, nonadiabatic holonomic gates are implemented by configuring the system’s Hamiltonian such that the time evolution operator parallel transports the computational space along a closed path, resulting in the desired gate as the holonomy. A Hamiltonian $H_{t}$ is said to be parallel transporting if its associated time evolution operator $U_{t}$ displaces the computational space in a parallel manner. An equivalent condition is that for any orthonormal basis $|v_{1}\rangle,|v_{2}\rangle,\dots,|v_{n}\rangle$ in the computational space,
@@
holds at all times [1, 2, 3, 20]. The parallel transport condition (3) ensures the Hamiltonian induces no couplings within the computational space.
@@ II.A   Gauge theoretic description of conventional nonadiabatic holonomic computation
An $n$-frame is a sequence of $n$ orthonormal vectors in $\mathcal{H}$. The set of all $n$-frames in $\mathcal{H}$ forms the Stiefel manifold $\mathcal{V}(n;\mathcal{H})$. In this paper, we will represent $n$-frames as row matrices, $V=(|v_{1}\rangle\,|v_{2}\rangle\dots|v_{n}\rangle)$. This allows us to apply standard matrix algebra to $n$-frames. In particular, we can act component-wise on an $n$-frame from the left with an operator, thereby generating a new row matrix of vectors that may or may not be a new $n$-frame. [^3] The resulting matrix of vectors is an $n$-frame if and only if the operator is an isometry on the span of the input frame. We can also act on an $n$-frame from the right with a real or complex matrix (with $n$ rows) and generate a new row matrix of vectors that are linear combinations of the original frame’s components; see [20] for details.
The Grassmann manifold $\mathcal{G}(n;\mathcal{H})$ is the manifold of $n$-dimensional subspaces of $\mathcal{H}$. Its topology and smooth structure are derived from its identification with the space of orthogonal projection operators on $\mathcal{H}$ of rank $n$, where each subspace is identified with the orthogonal projection onto that subspace.
Suppose the computational space $\mathcal{R}$ has dimension $n$, making it an element of $\mathcal{G}(n;\mathcal{H})$. A smooth transformation of $\mathcal{R}$ corresponds to a curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$, or equivalently to a curve of orthogonal projection operators $P_{t}$ of rank $n$, where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. The transformation is said to be closed if $\mathcal{R}_{\tau}=\mathcal{R}$.
@@
We say that a tangent vector of $\mathcal{V}(n;\mathcal{H})$ is horizontal if it is annihilated by $\mathcal{A}$, and we say that a curve in $\mathcal{V}(n;\mathcal{H})$ is horizontal if all its velocity vectors are horizontal. A fundamental result from the theory of fiber bundles asserts that for every smooth curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$ and every $n$-frame $V_{0}$ spanning $\mathcal{R}_{0}$, there is a unique horizontal curve $V_{t}$ in $\mathcal{V}(n;\mathcal{H})$ that starts at $V_{0}$ and is such that $V_{t}$ spans $\mathcal{R}_{t}$; see [21]. This curve is the horizontal lift of $\mathcal{R}_{t}$ starting at $V_{0}$. If $\mathcal{R}_{t}$ is a closed curve at $\mathcal{R}$, and $V_{t}$ is its horizontal lift, then both $V_{0}$ and $V_{\tau}$ span $\mathcal{R}$. The operators $\Pi_{t}=V_{t}V_{0}^{\dagger}$ are the parallel transport operators associated with $\mathcal{R}_{t}$, and the final operator $\Pi_{\tau}=V_{\tau}V_{0}^{\dagger}$ is the holonomy.
@@ II.B   The isoholonomic inequality and an execution time estimate for holonomic gates
where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. Suppose $\mathcal{R}_{t}$ is closed with holonomy $\Gamma$. The isoholonomic inequality for quantum gates states that the length of $\mathcal{R}_{t}$ is always greater than or equal to the isoholonomic bound
@@
From the isoholonomic inequality it follows that if a Hamiltonian $H_{t}$ generates the curve $\mathcal{R}_{t}$, meaning that $\mathcal{R}_{t}=U_{t}(\mathcal{R})$ where $U_{t}$ is the time-evolution operator associated with $H_{t}$, then the evolution time $\tau$ is not less than
$\tau[H_{t};\Gamma]=\frac{L(\Gamma)}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\rangle\!\rangle}.$ (8)
@@ II.C   Parallel transport in a rotating frame
Equation (3) describes the parallel transport condition relative to the laboratory frame. If the reference frame is changed, the parallel transport condition must be modified covariantly to maintain physical consistency. Here, we derive the equation that must be satisfied in a rotating reference frame to ensure that the parallel transport condition is satisfied in the laboratory frame.
@@
The potential $A_{t}=i\dot{R}_{t}R_{t}^{\dagger}$ compensates for the additional dynamics introduced by the time-varying reference frame.
Let $V$ be any computational frame, and write $|v_{j;t}\rangle$ for the evolution of the $j$th component of $V$. Then
$\langle v_{k;t}|\dot{v}_{l;t}\rangle=\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle,$ (13)
@@
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (14)
@@
$\langle v_{k;t}^{\textsc{rf}}|H^{\textsc{rf}}_{t}-A_{t}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (15)
@@
When changing the reference frame, care must also be taken to ensure that the computational space is transported in a loop, because if one of $\mathcal{R}_{t}$ and $\mathcal{R}_{t}^{\textsc{rf}}=R_{t}(\mathcal{R}_{t})$ forms a closed curve in $\mathcal{G}(n,\mathcal{H})$, the other may not necessarily do so. We have that $\mathcal{R}_{t}$ is a closed curve if and only if
$\mathcal{R}_{\tau}^{\textsc{rf}}=R_{\tau}R_{0}^{\dagger}(\mathcal{R}_{0}^{\textsc{rf}}).$ (16)
The computational space being transported in a loop can also be formulated more invariantly: Let $P$ be the orthogonal projection onto $\mathcal{R}$. That the computational space is transported in a loop over time $\tau$ can be expressed in the laboratory frame as $\langle\psi_{\tau}|P|\psi_{\tau}\rangle=1$ for all unit vectors $|\psi_{0}\rangle$ in $\mathcal{R}$. In the rotating frame picture, observables change with time according to Eq. (11), ensuring that all statistical quantities remain invariant under the change of reference frame. Consequently, the computational space is transported in a loop if and only if $\langle\psi_{\tau}^{\textsc{rf}}|P_{\tau}^{\textsc{rf}}|\psi_{\tau}^{\textsc{rf}}\rangle=1$ for all unit vectors $|\psi_{0}^{\textsc{rf}}\rangle$ in $\mathcal{R}_{0}^{\textsc{rf}}$.

### v3

@@ Abstract
Nonadiabatic holonomic quantum computation is a promising approach for implementing quantum gates that offers both efficiency and robustness against certain types of errors. A key element of this approach is a geometric constraint known as the parallel transport condition. According to the principle of covariance, this condition must be appropriately modified when changing reference frames. In this paper, we detail how to adjust the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame. Furthermore, building on gauge invariance considerations, we develop a framework for nonadiabatic holonomic quantum computation with projective gates. The parallel transport condition of this framework effectively addresses the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. We extend the isoholonomic inequality, which provides a fundamental bound on the efficiency of protocols used to implement holonomic quantum gates, to encompass projective quantum gates. We also determine a minimum execution time for projective holonomic quantum gates and show that this time can be attained when the codimension of the computational space is sufficiently large.
@@ I   Introduction
Nonadiabatic holonomic quantum computation [1, 2, 3] has been introduced as a fast alternative to its well-established, robust, but relatively slow adiabatic counterpart [4, 5, 6]. While adiabatic holonomic quantum computation relies on the gradual manipulation of control parameters, nonadiabatic holonomic quantum computation implements quantum gates through internal dynamics driven by parallel transporting Hamiltonians. These Hamiltonians, fundamental to the geometric nature of the gates, evolve the computational subspace without inducing any time-local rotation within it, ensuring that the computation is governed purely by geometric phases rather than dynamical effects.
In the laboratory frame, the parallel transport condition can be formulated as the requirement that the velocity of any computational basis remains within the kernel bundle of the non-Abelian Aharonov-Anandan connection [7, 8]. To simplify the analysis of quantum dynamical equations, a common approach is to switch to a rotating reference frame. However, to maintain physical consistency, the dynamical equations must be transformed covariantly when changing the reference frame. After outlining key concepts in holonomic quantum computing, we work out how to transform the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame.
@@
The isoholonomic inequality, a fundamental result in holonomic quantum computation [18, 19, 20, 21, 22], establishes a sharp lower bound on the length of the loops along which the computational subspace can be transported to implement a given quantum gate holonomically [21, 22]. We extend this inequality to projective quantum gates and, as a corollary, provide an estimate for the execution time of projective holonomic quantum gates. This extension offers deeper insight into the geometric constraints on the implementation of quantum gates in the projective setting. The paper concludes with a discussion of how to saturate the isoholonomic inequality and the execution time estimate.
@@
In this paper we use the word frame in two different senses. First, there are frames of reference that are used to describe the entire quantum system, called laboratory or rotating frames. These provide a basis for understanding the overall behavior of the system. Second, we introduce frames that span the computational space, which we call computational or $n$-frames. The context should make it clear which type of frame is being discussed.
@@ II   Conventional nonadiabatic holonomic quantum computation
Holonomic quantum computing uses holonomy to implement gates. Holonomy refers to the geometric phenomenon that when the computational space is transported along a closed path, a co-transported computational frame, that is, a frame that spans the computational space, may end up rotated relative to its original configuration, even though the transport is parallel, that is, performed without causing any time-local rotation of the computational frame.
A holonomic gate is a unitary operator representing the rotation of a computational frame resulting from a cyclic parallel transport of the computational space. A more precise description of a holonomic gate is as follows (see also [1, 2, 3, 21]). Consider a quantum system with a Hilbert space $\mathcal{H}$ and a computational subspace $\mathcal{R}$ [^1] Thus, computational input states are prepared to have support in $\mathcal{R}$ and are manipulated by unitary operators on $\mathcal{R}$, a.k.a gates. Let $\Pi_{t}$ be a one-parameter family of isometric embeddings of $\mathcal{R}$ in $\mathcal{H}$ that transport $\mathcal{R}$ in a loop:
@@
The transport is said to be parallel if, for each vector $|\psi\rangle$ in $\mathcal{R}$ and at all times $t$, the curve $|\psi_{t}\rangle=\Pi_{t}|\psi\rangle$ intersects the displaced computational space perpendicularly:
@@
Geometrically, condition (2) means that the operators $\Pi_{t}$ do not cause any time-local rotation within the computational space. The resulting rotation of a computational frame is thus solely a consequence of the translational motion of $\mathcal{R}$ through the Hilbert space. A crucial fact is that any curve of equidimensional subspaces of $\mathcal{H}$ can be realized as a transport by a unique one-parameter family of parallel transporting operators. These operators are the parallel transport operators associated with the curve. [^2] We assume that the parameter of any one-parameter family—a curve—of states, operators, or spaces ranges from 00 to $\tau$ and that the family depends piecewise smoothly on the parameter.
In practice, nonadiabatic holonomic gates are implemented by configuring the system’s Hamiltonian so that the time evolution operator parallel transports the computational space along a closed path with the desired gate as the holonomy. A Hamiltonian $H_{t}$ is said to be parallel transporting if its associated time evolution operator $U_{t}$ displaces the computational space in a parallel manner. An equivalent condition is that for any orthonormal basis $|v_{1}\rangle,|v_{2}\rangle,\dots,|v_{n}\rangle$ in the computational space,
@@
holds at all times [1, 2, 3, 21]. The parallel transport condition (3) ensures that the Hamiltonian induces no couplings within the computational space.
@@ II.A   Gauge theoretic description of conventional nonadiabatic holonomic computation
An $n$-frame is a sequence of $n$ orthonormal vectors in $\mathcal{H}$. The set of all $n$-frames in $\mathcal{H}$ forms the Stiefel manifold $\mathcal{V}(n;\mathcal{H})$. In this paper we will represent $n$-frames as row matrices, $V=(|v_{1}\rangle\,|v_{2}\rangle\dots|v_{n}\rangle)$. This allows us to apply standard matrix algebra to $n$-frames. In particular, we can act componentwise on an $n$-frame from the left with an operator, producing a new row matrix of vectors, which may or may not be a new $n$-frame. [^3] The resulting matrix of vectors is an $n$-frame if and only if the operator is an isometry on the span of the input frame. We can also act on an $n$-frame from the right with a real or complex matrix (with $n$ rows) and generate a new row matrix of vectors that are linear combinations of the components of the original frame; see [21] for details.
The Grassmann manifold $\mathcal{G}(n;\mathcal{H})$ is the manifold of $n$-dimensional subspaces of $\mathcal{H}$. Its topology and smooth structure are derived from its identification with the space of orthogonal projection operators on $\mathcal{H}$ of rank $n$, where each subspace is identified with the orthogonal projection onto that subspace. Suppose the computational space $\mathcal{R}$ has dimension $n$ and thus is an element of $\mathcal{G}(n;\mathcal{H})$. A smooth transformation of $\mathcal{R}$ corresponds to a curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$, or equivalently to a curve of orthogonal projection operators $P_{t}$ of rank $n$, where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. The transformation is said to be closed if $\mathcal{R}_{\tau}=\mathcal{R}$.
@@
We say that a tangent vector $X$ of $\mathcal{V}(n;\mathcal{H})$ is horizontal if $\mathcal{A}(X)=0$, and we say that a curve in $\mathcal{V}(n;\mathcal{H})$ is horizontal if all its velocity vectors are horizontal. A fundamental result from fiber bundle theory asserts that for every smooth curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$ and every $n$-frame $V_{0}$ spanning $\mathcal{R}_{0}$, there exists a unique horizontal curve $V_{t}$ in $\mathcal{V}(n;\mathcal{H})$ that starts at $V_{0}$ and is such that $V_{t}$ spans $\mathcal{R}_{t}$; see [23]. This curve is the horizontal lift of $\mathcal{R}_{t}$ starting at $V_{0}$. If $\mathcal{R}_{t}$ is a closed curve at $\mathcal{R}$ and $V_{t}$ is its horizontal lift, then both $V_{0}$ and $V_{\tau}$ span $\mathcal{R}$. The family of operators $\Pi_{t}=V_{t}V_{0}^{\dagger}$ are the parallel transport operators associated with $\mathcal{R}_{t}$, and the final member of the family $\Pi_{\tau}=V_{\tau}V_{0}^{\dagger}$ is the holonomy. The parallel transport operators transport any frame $V$ in $\mathcal{R}$ according to $V_{t}=\Pi_{t}V$, and the transformation from $V$ to $V_{\tau}$ is described by the holonomy, $V_{\tau}=\Pi_{\tau}V$.
@@ II.B   The isoholonomic inequality and an execution time estimate for holonomic gates
where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. Suppose $\mathcal{R}_{t}$ is closed with the holonomy $\Gamma$. The isoholonomic inequality for quantum gates says that the length of $\mathcal{R}_{t}$ is always greater than or equal to the isoholonomic bound of $\Gamma$,
@@
From the isoholonomic inequality it follows that if a Hamiltonian $H_{t}$ generates the curve $\mathcal{R}_{t}$, meaning that $\mathcal{R}_{t}=U_{t}(\mathcal{R})$, where $U_{t}$ is the time evolution operator associated with $H_{t}$, then the evolution time $\tau$ is not less than
$\tau_{\textsc{qsl}}[H_{t};\Gamma]=\frac{L(\Gamma)}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\rangle\!\rangle}.$ (8)
@@
$\tau\geq\tau_{\textsc{qsl}}[H_{t};\Gamma]$ (10)
holds in particular for parallel transporting Hamiltonians, $\tau_{\textsc{qsl}}[H_{t};\Gamma]$ serves as a lower bound—a quantum speed limit—on the time it takes to execute $\Gamma$ holonomically.
@@ II.C   Parallel transport in a rotating frame
Equation (3) describes the parallel transport condition relative to the laboratory frame. When the reference frame is changed, the parallel transport condition must be modified covariantly to maintain physical consistency. Here, we derive the equation that must be satisfied in a rotating reference frame to ensure that the parallel transport condition is satisfied in the laboratory frame.
@@
The potential $A_{t}^{\textsc{rf}}=i\dot{R}_{t}R_{t}^{\dagger}$ compensates for the additional dynamics introduced by the time-varying reference frame.
Let $V$ be any computational frame, and write $|v_{k;t}\rangle$ for the evolution of the $k$th component of $V$. Then
$\langle v_{k;t}|\dot{v}_{l;t}\rangle=\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}^{\textsc{rf}}|v_{l;t}^{\textsc{rf}}\rangle,$ (14)
@@
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}^{\textsc{rf}}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (15)
@@
$\langle v_{k;t}^{\textsc{rf}}|H^{\textsc{rf}}_{t}-A_{t}^{\textsc{rf}}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (16)
@@
When working in a rotating frame, it is important to check whether the computational space is transported in a loop in the laboratory frame picture. Suppose $\mathcal{R}^{\textsc{rf}}_{t}$ is the trajectory of the computational space in the rotating frame picture. The trajectory of the computational space $\mathcal{R}$ in the laboratory frame is then given by $\mathcal{R}_{t}=R_{t}^{\dagger}(\mathcal{R}_{t}^{\textsc{rf}})$, which is closed if and only if $R_{\tau}^{\dagger}(\mathcal{R}_{\tau}^{\textsc{rf}})=R_{0}^{\dagger}(\mathcal{R}^{\textsc{rf}})$. This may require that $\mathcal{R}^{\textsc{rf}}_{t}$ is not closed. Figure 1 illustrates a scenario where the computational space follows a closed curve in the laboratory but not in the rotating frame picture.
[…]
Figure 1: The left figure illustrates a scenario where the computational space $\mathcal{R}$ is transported along a closed curve $\mathcal{R}_{t}$ in the laboratory frame. In contrast, the right figure shows that this does not necessarily imply that the corresponding curve in the rotating frame, $\mathcal{R}_{t}^{\textsc{rf}}$, is also closed. This discrepancy arises when the initial and final rotation operators, $R_{0}$ and $R_{\tau}$, map $\mathcal{R}$ into different subspaces. Similarly, if the computational space is transported along a closed curve in the rotating frame, it does not necessarily follow that it traces a closed curve in the laboratory frame.
The computational space being transported in a loop can also be formulated more invariantly: Let $P$ be the orthogonal projection onto $\mathcal{R}$. That $\mathcal{R}_{t}$ is closed can be expressed as $\langle\psi_{\tau}|P|\psi_{\tau}\rangle=1$ for all unit vectors $|\psi\rangle$ in $\mathcal{R}$. In the rotating frame picture, observables evolve over time according to Eq. (12), ensuring that all statistical quantities remain invariant under the change of reference frame. Consequently, $\mathcal{R}_{t}$ is closed if and only if $\langle\psi_{\tau}^{\textsc{rf}}|P_{\tau}^{\textsc{rf}}|\psi_{\tau}^{\textsc{rf}}\rangle=1$ for all unit vectors $|\psi^{\textsc{rf}}\rangle$ in $\mathcal{R}^{\textsc{rf}}$.